# Voting by mail: a Markov chain model for managing the security risks of absentee voting

### v2

@@ 6.2 Malicious attack scenarios given baseline
Next, we introduce malicious attacks to the baseline. We focus on malicious attacks and . Malicious attacks are modeled to last a single day during the 49-day time horizon to reflect feasible real-world election interference events, and we vary the attack strength of each attack to consider different scenarios. Malicious attacks therefore, modify the transition probability matrix () for a single day associated with the attack . This approach allows us to assess the impact of attack timings and the mitigation strategies for countering attacks.
@@ 6.2.1 X9: Malicious attack to challenge the signature of a valid ballot
Malicious attack captures bad actors in an election office erroneously rejecting marked ballots by challenging their signatures. This scenario, with a base arc transition probability of 0.055, is set to affect roughly one of 19 precincts in Milwaukee. Referencing Figure 2, malicious attack impacts process state , where ballots are collected and verified by the election office. On a specific day, this could lead to a significant number of ballots being improperly discarded.
@@ 6.2.2 X13: Malicious attack discard ballot in transit via USPS
Malicious attack represents bad actors in a postal office who discard completed ballots at the sorting or storage point. Referring back to Figure 2, we see that malicious attack directly impacts process state where the voter returns the completed ballot via USPS. In Milwaukee County, a large-scale attack affects one of the 36 USPS offices for a day [13]. Consequently, there is a probability of 0.033 that ballots in process state are impacted by the attack.
@@ 6.2.3 X29: Masquerade Attacks, Ballot Modifying
We then consider the malicious attack , which alters voters’ original ballots and directly impacts process states . The consequence of this attack leads to altered ballots that are both counted and uncounted (final ballot states (C,A) and (NC,A)). The arc probability strengths from attack are reflected here, since there exists no known reference of an executed masquerade attack.
We assume that there is an equal probability that the altered ballot are counted or uncounted. The analysis regarding attack underscores the vulnerability of the voting process to pre-altered ballot distributions. This manipulation targets the early stages of ballot circulation, redirecting ballots from the process layer state to the attack layer state .
@@ 6.2.4 Attack timing
Figure 5 illustrates the deviations in the expected number of (C,U) ballots associated with each attack scenario as a function of the day of the attack. All values are negative, indicating that each attack reduces the number of ballots in the preferred (C,U) state. Attack , which models the false rejection of valid ballots at the election office (process state VII), experiences its largest deviation in (C,U) ballots when the attack is initiated on day 38. This attack peak aligns with a surge of ballots routed to the election office via early voting (M6), which begins on day 36 and introduces a direct arc from process state II to state VII, in addition to those being returned by mail or drop box. Similarly, attack , representing the loss of ballots in USPS transit, peaks on day 37 in terms of its deviation in the expected number of (C,U) ballots. This deviation increases due to ballot return reminders (M5) that start to alert voters on Day 36. This results in an increase in the flow of ballots in transit, which results in more ballots that are available to be maliciously stolen.
Attack , a masquerade attack involving stolen or pre-marked ballots (state II), experiences its largest deviation in (C,U) ballots earlier in the election cycle on day 7, which results in 706 fewer (C,U) ballots. This occurs due to the initial surge in ballot issuance and early returns at this time. A secondary deviation of 614 (C,U) ballots occurs on day 36, coinciding with an increase in voters returning ballots due to automatic ballot reminders (M5). These results indicate that attack timing–—specifically its alignment with the flow of ballots–—is crucial for determining the consequences of an attack. The consequences of an attack corresponds to its interaction with process dynamics during periods of high ballot flow.
@@ 6.3 Worst-Case Scenario Modeling
We build on the attack scenarios in Section 6.2.4 to create a “worst-case scenario” to evaluate various policy implementations in the VBM process under severe conditions as well as the impact on mitigations. The worst-case scenario involves the simultaneous execution of three malicious attacks within the 49-day election cycle. Figure 5 illustrates the highest impact for each attack as: on day 38, on day 37, and on day 7. The following arc probabilities are all set to 0.10 on these days: arc from node to on day 37; arc from node to on day 7; arc from node to . We also consider a sensitivity analysis that evaluates the impact of mitigation strength and associated strength against the worst-case attack scenario. Referencing Table 3 in Section 3.4, four mitigations counter the attacks in the worst-case scenario (see Table 3):
@@
Next, we examine varying mitigations automatic ballot notifications and replacement ballots simultaneously to reflect their interdependency. We find that deploying automatic ballot notifications with the same availability as replacement ballots, at , results in 325,584 counted (C,U) ballots, an increase of 9,742 ballots as compared to the baseline, worst-case performance in the first row of Table 11. An automatic ballot notifications availability of 0.90 results in 62 expected (C,A) ballots. Lastly, since automatic ballot notifications and early voting can function interdependently, we examine the impact of simultaneously varying automatic ballot notifications and early voting . When deployed at 0.90 availability, we observe 325,485 (C,U) ballots, similar to the previous observations. Additionally, we observe fewer ballots not returned, with 17,243 (NC,NR) ballots compared with 20,371 (NC,NR) ballots in the Milwaukee baseline without malicious attacks. We conclude that in-person absentee voting gives voters a powerful recourse option to counter malicious and non-malicious attacks.
@@
Next, we study the sensitivity of the DTMC model mitigation strength. The two final ballot states we study are the desired final ballot state (C,U) and the unreturned ballots (NC,NR) since these two final ballot states are the most impacted by mitigations. However, other ballot states are impacted by mitigations to a lesser degree. We perform a one-way sensitivity analysis by changing each mitigation strength by from its baseline value (see Table 2). Figure 6 shows a tornado graph illustration of the results. The blue half of Figure 6 on the left represents the difference in the number of correctly counted and unaltered ballots when the mitigation strength decreases. The red half is the difference in the number of ballots that are not counted and not returned, which occurs when mitigation strength increases. The tornado graph shows that the DTMC model is most sensitive to drop boxes and automatic ballot notifications . In contrast, replacement ballots and early voting (independently) do not significantly affect the outputs in a significant manner. However, their impact can be larger when their availability is considered within the context of other mitigations (see Table 11). These results demonstrate that modeling attack timing in conjunction with mitigation availability provides crucial insight into the consequence of each attack.
@@ 7 Conclusions
VBM process and policy are rapidly evolving in the United States. Our research presents a quantitative tool designed to assess the risks and impacts of mitigations within VBM systems. We introduce a discrete-time Markov Chain modeling framework, which we apply to a case study based on the 2020 General Election in Milwaukee County to evaluate the security and effectiveness of VBM processes. Our findings highlight the critical role of dynamic mathematical modeling in computational risk analysis for VBM systems. This layered approach effectively organizes the complex interactions among cyber, physical, and human elements in VBM, providing insight into VBM vulnerabilities. Our analysis, which leverages various ballot outcomes and real-world data, highlights the model’s ability to assess voting risks.
Singular mitigations, such as drop boxes, can bolster the VBM system, while other mitigations must be combined to reduce risk. A crucial observation is that automatic ballot notifications (dependent on replacement ballots or early voting) is the only mitigation capable of effectively countering ballot-altering malicious attacks. The variation of each mitigation policy set changes the VBM process’s resilience to the worst-case scenario. These changes mark quantifiable metrics for evaluating VBM security and highlight the need for an integrated and adaptive mitigation approach to ensure the integrity of the VBM process.
@@
There are several limitations of our study that provide insight into topics for future research. First, this analysis considers malicious attacks that occur on a single-day and is limited by a lack of historical data regarding malicious attacks. Future research could diversify attack scenarios and incorporate a broader range of data to enhance the model’s sensitivity to resource constraints and varying electoral processes. Second, future research could account for resource constraints and system congestion. Lastly, we did not examine the likelihood of attack scenarios altering election outcomes. Successful attacks on voting systems compromise the integrity of elections and undermine public trust in political systems, regardless of whether they change the outcomes. Although not all attacks aim to change election results, this remains a significant concern that future research could investigate.
@@ Acknowledgments
This work was in part funded by the National Science Foundation Award 2000986. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the National Science Foundation.

### v3

@@ 6.2 Malicious attack scenarios given baseline
Next, we introduce malicious attacks to the baseline. We focus on malicious attacks and to illustrate the three attack types—fail rate increase, ballot modifying, and process altering. Malicious attacks are modeled to last a single day during the 49-day time horizon to reflect feasible real-world election interference events, and we vary the attack strength of each attack to consider different scenarios. We consider worst-case attack strengths to stress-test the model. Malicious attacks therefore, modify the transition matrix () for a single day associated with the attack . This approach allows us to assess the impact of attack timings and the mitigation strategies for countering attacks.
@@ 6.2.1 X9: Malicious attack to challenge the signature of a valid ballot
Malicious attack captures bad actors in an election office erroneously rejecting marked ballots by challenging their signatures. This scenario, with a base arc transition of 0.055, is set to affect roughly one of 19 precincts in Milwaukee. Referencing Figure 2, malicious attack impacts process state , where ballots are collected and verified by the election office. On a specific day, this could lead to a significant number of ballots being improperly discarded.
@@ 6.2.2 X13: Malicious attack discard ballot in transit via USPS
Malicious attack represents bad actors in a postal office who discard completed ballots at the sorting or storage point. Referring back to Figure 2, we see that malicious attack directly impacts process state where the voter returns the completed ballot via USPS. In Milwaukee County, a large-scale attack affects one of the 36 USPS offices for a day [15]. Consequently, there is a relative likelihood of 0.033 that ballots in process state are impacted by the attack.
@@ 6.2.3 X29: Masquerade Attacks, Ballot Modifying
We then consider the malicious attack , which alters voters’ original ballots and directly impacts process states . The consequence of this attack leads to altered ballots that are both counted and uncounted (final ballot states (C,A) and (NC,A)). Since there exists no known reference of an executed masquerade attack, we assume that there is an equal likelihood that the altered ballot are counted or uncounted. The manipulation from malicious attack targets the early stages of ballot circulation, redirecting ballots from the process layer state to the attack layer state .
@@ 6.2.4 Attack timing
Figure 5 illustrates the deviations in the expected number of (C,U) ballots associated with each attack scenario as a function of the day of the attack. All values are negative, indicating that each attack reduces the number of ballots in the preferred (C,U) state. Attack , which models the false rejection of valid ballots at the election office (process state VII), experiences its largest deviation in (C,U) ballots when the attack is initiated on day 38. This attack peak aligns with a surge of ballots routed to the election office by mail via early voting (M6), which begins on day 36 and introduces a direct arc from process state II to state VII, in addition to those being returned by mail or drop box. Similarly, attack , representing the loss of ballots in USPS transit, peaks on day 37 in terms of its deviation in the expected number of (C,U) ballots. This is due to ballot return reminders (M5) beginning on Day 36. This prompts voters to return their ballots and results in an increase in the flow of ballots in transit, which results in more ballots being available to be maliciously stolen by malicious actors. However, attack affects few ballots in the week prior to election day, when most ballots have already been returned.
Attack , a masquerade attack involving stolen or pre-marked ballots (state II), experiences its largest deviation in (C,U) ballots earlier in the election cycle on day 7, which results in 706 fewer (C,U) ballots. This occurs due to the initial surge in ballot issuance and early returns at this time. A secondary deviation of 614 (C,U) ballots occurs on day 36, coinciding with an increase in voters returning ballots due to automatic ballot reminders (M5). These results indicate that attack timing–—specifically its alignment with the flow of ballots–—is crucial for determining the consequences of an attack. The consequences of an attack corresponds to its interaction with process dynamics during periods of high ballot flow. As with , affects few ballots in the week prior to election day.
@@ 6.3 Worst-Case Scenario Modeling
We build on the attack scenarios in Section 6.2.4 to create a “worst-case scenario” to evaluate various policy implementations in the VBM process under severe conditions as well as the impact on mitigations. The findings presented in this section are contingent upon the assumed strengths of these mitigations and the strengths of the associated attacks.
The worst-case scenario involves the simultaneous execution of three malicious attacks within the 49-day election cycle. Figure 5 illustrates the highest impact for each attack as on day 38, on day 37, and on day 7. The following arc transitions are all set to 0.10 on these days: arc from node to on day 37; arc from node to on day 7; arc from node to . We also consider a sensitivity analysis that evaluates the impact of mitigation strength and associated strength against the worst-case attack scenario. Referencing Table 3 in Section 3.4, four mitigations counter the attacks in the worst-case scenario (see Table 3):
@@
Next, we examine varying mitigations automatic ballot notifications and replacement ballots simultaneously to reflect their interdependency. We find that deploying automatic ballot notifications with the same availability as replacement ballots, at , results in 325,584 counted (C,U) ballots, an increase of 9,742 ballots as compared to the baseline, worst-case performance in the first row of Table 11. An automatic ballot notifications availability of 0.90 results in 62 expected (C,A) ballots. Lastly, since automatic ballot notifications and early voting can function interdependently, we examine the impact of simultaneously varying automatic ballot notifications and early voting . When deployed at 0.90 availability, we observe 325,485 (C,U) ballots, similar to the previous observations. Additionally, we observe fewer ballots not returned, with 17,243 (NC,NR) ballots compared with 20,371 (NC,NR) ballots in the Milwaukee baseline without malicious attacks. We conclude that early absentee voting gives voters a powerful recourse option to counter malicious and non-malicious attacks.
@@
Next, we study the sensitivity of the DTMC model mitigation strength. The two final ballot states we study are the desired final ballot state (C,U) and the unreturned ballots (NC,NR) since these two final ballot states are the most impacted by mitigations. However, other ballot states are impacted by mitigations to a lesser degree. We perform a one-way sensitivity analysis by changing each mitigation strength by from its baseline value (see Table 2). Figure 6 shows a tornado graph illustration of the results. The blue half of Figure 6 on the left represents the difference in the number of correctly counted and unaltered ballots when the mitigation strength decreases. The red half is the difference in the number of ballots that are not counted and not returned, which occurs when mitigation strength increases. The tornado graph shows that the DTMC model is most sensitive to drop boxes and automatic ballot notifications . In contrast, replacement ballots and early voting (independently) do not significantly affect the outputs in a significant manner. However, their impact can be larger when their availability is considered within the context of other mitigations (see Table 11). These results demonstrate that modeling attack timing in conjunction with mitigation availability provides crucial insight into the consequence of each attack. Our findings depend on the assumed strengths of these mitigations and the strengths of the associated attacks. The number of affected ballots in real settings would also depend on voter behavior and the operational reliability of election processes.
@@ 7 Conclusions
VBM process and policy are rapidly evolving in the United States. Our research presents a quantitative tool designed to stress-test Vote-by-Mail systems. We introduce a discrete-time Markov Chain modeling framework, which we apply to a case study based on the 2020 General Election in Milwaukee County to evaluate the security and effectiveness of VBM processes. The case study analysis frames threats as hypothetical scenarios to compare the relative impact of defensive mitigations. Considering “worst-case” conditions with threat actors allows us to identify which mitigations have the greatest effect, independent of the actual (and historically low) base rate of attack. Our findings highlight the critical role of dynamic mathematical modeling in computational risk analysis for VBM systems. This layered approach effectively organizes the complex interactions among cyber, physical, and human elements in VBM, providing insight into VBM vulnerabilities. Our analysis, which leverages various ballot outcomes and real-world data, is framed as a stress-test of systemic vulnerabilities rather than a literal forecast of election outcomes. By evaluating various hypothetical, worst-case scenarios, the model serves to identify aspects of the VBM infrastructure that are most sensitive to disruption.
Our analysis suggests that singular mitigations, such as drop boxes, can bolster the VBM system, while other mitigations must be combined to reduce risk. A crucial observation is that automatic ballot notifications (dependent on replacement ballots or early voting) is the only mitigation capable of effectively countering ballot-altering malicious attacks. The variation of each mitigation policy set changes the VBM process’s resilience to the worst-case scenario. These changes mark quantifiable metrics for evaluating VBM security and highlight the need for an integrated and adaptive mitigation approach to ensure the integrity of the VBM process.
@@
There are several limitations of our study that provide insight into topics for future research. First, this analysis considers malicious attacks that occur on a single-day and is limited by a lack of historical data regarding malicious attacks. Future research could diversify attack scenarios and incorporate a broader range of data to enhance the model’s sensitivity to resource constraints and varying electoral processes. Second, future research could account for resource constraints, system congestion, and additional mitigations such as enhanced voter education and ballot design. Third, we did not examine the likelihood of attack scenarios altering election outcomes. Successful attacks on voting systems compromise the integrity of elections and undermine public trust in political systems, regardless of whether they change the outcomes. Although not all attacks aim to change election results, this remains a significant concern that future research could investigate.
Finally, there is significant epistemic uncertainty in this application area. It assumes that the defined states and transitions capture the full spectrum of possible ballot movements and threats. However, this structure cannot inherently predict “unknown unknowns,” such as entirely new attack vectors or rapid legislative shifts that fundamentally alter the voting process. The results presented in this paper should be interpreted as a risk assessment of the current security landscape. The model is designed to be extensible; therefore, the underlying Markov chain structure must be updated as new threats or process change emerge.
@@ Acknowledgments
This work was in part funded by the National Science Foundation Award 2000986. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the National Science Foundation. The authors would like to thank the anonymous reviewers and the associate editor, whose suggestions for improvement led to a substantially improved manuscript.